OpenStudy (nsh4267):

Find the solutions of the equation. x^3+3x^2+4x=0

OpenStudy (anonymous):

well just factor, you can simplify by taking out an x: x ( x^2 + 3x - 4) and then factor that into: x (x+4)(x-1) and then the solution's include 0, -4, and 1

OpenStudy (nsh4267):

that is not what the back of the book says

OpenStudy (anonymous):

what does the back of the book say.

OpenStudy (nsh4267):

\[0, -\frac{ 3 }{ 2 }\pm \frac{ 1 }{ 2 }, \sqrt{7i} \]

OpenStudy (anonymous):

type your equation again please cause the equation you typed has the answers i provided.

OpenStudy (nsh4267):

\[x^{3}+3x^{2}+4x=0\]

OpenStudy (anonymous):

oh ok..sorry so you can still factor out the x leaving you with (x^2 + 3x+4) and then use the quadratic formula..do you know the quadratic formula?

OpenStudy (nsh4267):

yes okay. now I know how they got it